# BoostStep: Boosting mathematical capability of Large Language Models via improved single-step reasoning

### v2

@@ 4.1 Experiment setting
Reasoning Model.  Our primary reasoning model is GPT-4o Hurst et al. (2024). To demonstrate the generality of our approach, we also conducted tests on Qwen2.5-Math-72B-Instruct Yang et al. (2024), which is the state-of-the-art open-source model in mathematical reasoning capabilities.
Evaluation Benchmark.  We tested our approach on several challenging open-source mathematical benchmarks, including MATH500 Hendrycks et al. (2021), AQuA Ling et al. (2017), OlympiadBench-TO He et al. (2024) and MATHBench Liu et al. (2024a) College-level and High-level tasks. In addition, we manually collected a selection of problems from the AMC-10 and AMC-12 competitions to serve as even more challenging benchmarks. * The AMC test set is available at https://github.com/beichenzbc/BoostStepTo simulate benchmarks with lower similarity to the example problem bank, we also conducted tests on MathVision Wang et al. (2024) and MathVerse Zhang et al. (2025), highly challenging multi-modal math benchmarks
@@
Retriever.  We utilized the classic TF-IDF encoding method combined with cosine similarity as the retriever for all for both problem-level and step-level in-context learning methods. The TF-IDF weight matrix is derived from the example problem bank. This is because the impact of the newly generated step on both TF and IDF is negligible, and real-time calculation of TF-IDF would require a significant amount of time.
@@ 4.2 Comparing to Problem-Level ICL
Model Method in-domain out-domain
MATH AMC12 AMC10 AQUA MathBench(C) MathBench(H) OlympiadBench-TO
GPT-4o 0-shot 73.4 53.6 55.8 81.1 80.0 77.3 40.6
few-shot 73.8 56.5 56.7 83.9 80.7 79.3 39.3
Ours 76.4 63.0 60.4 85.4 82.0 84.0 43.3
Qwen 0-shot 83.0 67.4 67.7 84.6 80.6 82.0 49.7
few-shot 83.8 67.4 66.8 85.0 81.3 82.7 49.9
Ours 85.2 69.2 69.6 86.6 82.7 84.7 52.7
@@
Better Performance. We compare the performance between traditional few-shot learning and our step-level in-context learning across multiple benchmarks and base models. The results are presented in Tab. 1. Our step-level in-context learning achieves a general and significant improvement across various benchmarks compared to problem-level few-shot learning.
@@
Method MathVision-Mini MathVerse-Mini
0-shot 30.6 53.2
few-shot 28.7 (-1.9) 53.2 (0.0)
Ours 35.2 (+4.6) 54.2 (+1.0)
Table 3: Experiments on the sensitivity of the similarity between the question and the example problem bank. R_t indicates that the examples are the t_th similar for different methods without any rejection strategy. Given a less similar example, our method suffers a 2.26% performance loss, which is much lower compared to few-shot learning(4.4%).
@@
few-shot R_1 52.2 56.5 56.7
few-shot R_4 46.3 (-5.9) 52.2(-4.3) 53.7 (-3.0)
Ours R_1 56 62.3 60.4
Ours R_4 52.2 (-3.8) 61.6 (-0.7) 58.1 (-2.3)
Better Generalizability. Traditional few-shot learning requires the example problem bank highly similar to the questions to be solved, which limits its generalizability. To compare the generalizability, We also test different methods on multi-modal mathematical benchmarks including MathVision Wang et al. (2024) and MathVerse Zhang et al. (2025), which has much lower similarity with our example problem bank. The results are shown in Tab. 2. Problem-level few-shot learning not only fails to enhance reasoning performance but can also have a negative impact, while our method continues to achieve appreciable improvements, demonstrating better general applicability.
Better Robustness. We also manually decrease the similarity between the examples and the problems by selecting the t_th similar example during reasoning to evaluate the robustness. The result is shown in Tab. 3. We can observe that traditional problem-level in-context learning suffers from a severe decrease and is even worse than 0-shot learning when t is larger than 4. In contrast, our method does not show a significant decline and is consistently better than the 0-shot reasoning.
@@ 4.3 Construction of Example Problem Bank
To better align with the steps in reasoning, we propose constructing a step-level question bank based on the reasoning content rather than grammatical divisions. To prove our assumption, we compare our approach with a commonly used strategy that constructs steps based on grammatical segmentation, using periods ’.’ as the delimiter, on the same dataset PRM800K and under identical conditions. Results are presented in Tab. 4. Our method largely outperforms those using periods as a delimiter.
@@ 4.4 Comparison of Retrieving Strategies
The key factor in determining the effectiveness of in-context learning lies in the relevance of the retrieved examples. At the finer-grained step level, designing an appropriate retrieval strategy becomes even more crucial and challenging. Therefore, we propose the first-try strategy, which involves understanding what the model currently needs to reason about using a first attempt and then searching the problem set for similar steps to guide the model in fully outputting the current step. To validate the effectiveness of this method, we compare it with several other strategies mentioned in Sec.3.3, retrieving by the entire reasoning path $s_{i-1},s_{i-2},\ldots,s_{1},q$ or only by the immediately preceding step $s_{i-1}$.
@@
Table 5: Comparison of different retrieval strategies in step-level in-context learning. The base model is GPT-4o and all the prompts are the same. ’Path’ represents retrieving by the reasoning path including all previous steps $s_{i-1},s_{i-2},\ldots,s_{1}$ and question $q$, while ’Pre-Step’ represents retrieving by only the immediately preceding step $s_{i-1}$. The best results are in bold.
@@ 4.5 Example-guided Step-level Tree Search
Reason Verify AMC12 AMC10 MATH
w/o Tree-Search 53.6 55.8 73.4
58.7 59.0 77.8
64.4 62.2 79.2
61.6 60.4 78.2
65.2 63.6 79.4
As discussed above, the reasoning capability of the generating model and the verifying capability of the critic model are two core factors of Step-level Tree Search methods, and our strategy can bring benefits in both ways. On one hand, it can improve the accuracy of generating candidate nodes using the previously mentioned first-try strategy when reasoning nodes are generated. On the other hand, it can increase the accuracy of evaluation by introducing similar examples during critic model assessments and therefore ensures that the correct reasoning nodes are more likely to be preserved. These two strategies can be decoupled, allowing us to demonstrate the effectiveness of each component through ablation studies.
In the setup for tree search methods, we utilize GPT-4o as the reasoning model and GPT-4o-mini as the Process-supervised Reward Model. Moreover, we adopted the Pairwise Preference Reward Model (PPRM) configuration Zhang et al. (2024b)which reduces variability and ensures a more robust evaluation. Detailed settings will be listed in the appendix.
Tab. 6 presents the results of integrating in-context learning into the reasoning and evaluation phases of Tree Search methods. The results indicate that introducing examples enhances the verifying and reasoning capabilities of tree search methods, therefore both approaches contribute to the improvement of overall reasoning performance by providing similar ground truth steps. Specifically, introducing example steps in reasoning phases will bring more benefits to the overall reasoning quality.
4.6 Case Study
@@ 5 Conclusion
We propose step-level in-context learning, which provides real-time, fine-grained guidance during the reasoning process by searching for similar steps according to the first-try reasoning attempt. This approach improves the model’s reasoning capabilities and reduces the dependency on the similarity of the example problem set, thereby increasing the generalizability of in-context learning. Moreover, our method can also enhance the reasoning and evaluation capability of Step-level Tree Search methods by introducing similar steps in reasoning and verifying phases respectively, thereby improving the overall reasoning correctness.
@@ A.3 Benchmarks
For multi-modal math benchmarks, we use MathVision-Mini Wang et al. (2024) and vision-dominant version of problems in MathVerse-Mini  Zhang et al. (2025).
Appendix B Detailed Setup for Step-Level Tree Search
In the setup for tree search methods, we utilize GPT-4o as the reasoning model and employ GPT-4o-mini as the Process-supervised Reward Model (PRM). For the PRM, we adopted the Pairwise Preference Reward Model (PPRM) configuration Zhang et al. (2024b). Specifically, PPRM transforms the absolute rewards calculation into preference predictions between solutions to calculate rewards. This approach reduces the variability associated with scoring characteristics and thus leads to a more robust and consistent evaluation of different solutions.

### v3

@@ 4.1 Experiment setting
Reasoning Model.  Our primary reasoning model is GPT-4o Hurst et al. (2024). To demonstrate the generality, we also conducted tests on Qwen2.5-Math-72B-Instruct Yang et al. (2024). Moreover, current SOTA reasoning models Qwen-QwQ-32B Team (2024) and DeepSeek-R1-671B Guo et al. (2025) were also included in our experiment.
Evaluation Benchmark.  We tested our approach on several challenging open-source mathematical benchmarks, including MATH500 Hendrycks et al. (2021), AQuA Ling et al. (2017), OlympiadBench-TO He et al. (2024) and MATHBench Liu et al. (2024a) College-level and High-level tasks. In addition, we manually collected a selection of problems from the AMC-10 and AMC-12 competitions to serve as even more challenging benchmarks. 222 * The AMC test set is available at https://github.com/beichenzbc/BoostStep To simulate benchmarks with lower similarity to the example problem bank, we also conducted tests on MathVision Wang et al. (2024a) and MathVerse Zhang et al. (2025), highly challenging multi-modal math benchmarks
@@
Retriever.  We utilized the classic TF-IDF encoding method combined with cosine similarity as the retriever for all methods. The TF-IDF weight matrix is derived from the example problem bank because the impact of the newly generated step is negligible, and real-time calculation of TF-IDF would require a significant amount of time.
@@ 4.2 Comparing to Problem-Level ICL
Model Method in-domain out-domain Avg
MATH AMC12 AMC10 AQUA MathBench(C) MathBench(H) OlympiadBench
GPT-4o 0-shot 73.4 53.6 55.8 81.1 80.0 77.3 40.6 66.0
few-shot 73.8 56.5 56.7 83.9 80.7 79.3 39.3 67.2 (+1.2)
Ours 76.4 63.0 60.4 85.4 82.0 84.0 43.3 70.6 (+4.6)
Qwen 0-shot 83.0 67.4 67.7 84.6 80.6 82.0 49.7 73.6
few-shot 83.8 67.4 66.8 85.0 81.3 82.7 49.9 73.8 (+0.2)
Ours 85.2 69.2 69.6 86.6 82.7 84.7 52.7 75.8 (+2.2)
@@
Performance We compare the performance between traditional few-shot learning and our step-level in-context learning across multiple benchmarks and base models. The results are presented in Tab. 1. Our step-level in-context learning achieves a general and significant improvement across various benchmarks compared to problem-level few-shot learning.
@@
Method MathVision MathVerse Avg
0-shot 30.6 53.2 41.9
few-shot 28.7 (-1.9) 53.2 (0.0) 41.0 (-0.9)
Ours 35.2 (+4.6) 54.2 (+1.0) 44.7 (+2.8)
Table 3: Experiments on the sensitivity of the similarity between the question and the example problem bank. R_t indicates that the examples are the t_th similar without any rejection strategy.
@@
few-shot R_1 52.2 (+1.5) 56.5 (+2.9) 56.7 (+0.9)
few-shot R_4 46.3 (-4.4) 52.2 (-1.4) 53.7 (-2.1)
Ours R_1 56.0 (+5.3) 62.3 (+8.7) 60.4 (+4.6)
Ours R_4 52.2 (+1.5) 61.6 (+8.0) 58.1 (+2.3)
Potential A key focus of in-context learning is determining how difficult a particular example can effectively guide the new problems, indicating the potential of these methods. Problem-level in-context learning faces significant challenges in leveraging simpler examples to enhance the model’s reasoning performance on more difficult questions. However, our strategy offers guidance at the step level, thereby overcoming this upper limit. To validate this, we select SOTA reasoning models QwQ-32B-Preview Team (2024) and DeepSeek R1 Guo et al. (2025) and utilized simpler example problems from PRM800K to guide the reasoning on the most challenging mathematical benchmark AIME MAA (2024). The results are shown in tab. 4, which indicate that traditional few-shot learning fails to provide effective guidance while our strategy demonstrates continuous improvement, demonstrating that it can boost the most advanced reasoning models on the most challenging tasks with a much simpler example.
Table 4: Experiment on "simple-aids-difficult" potential. We use simpler example problems from PRM800K to guide SOTA reasoning models Deepseek-R1 and Qwen-QwQ on the most challenging mathematical benchmarks AMC 12 and AIME. Considering that the AIME consists of only 30 questions each year, making the results prone to fluctuations, we evaluated the questions three times annually and reported the average accuracy.
Model Method AMC12 AIME23 AIME24
QwQ 0-shot 79.7 38.9 43.3
few-shot 81.2 33.3 (-5.6) 38.9 (-4.4)
Ours 88.4 41.1 (+2.2) 47.8 (+4.5)
DS-R1 0-shot 94.2 75.6 80.0
few-shot 97.1 65.6 (-10.0) 70.0 (-10.0)
Ours 97.1 77.8 (+2.2) 82.2 (+2.2)
Generalizability Traditional few-shot learning requires the example problem bank highly similar to the questions to be solved, which limits its generalizability. To compare the generalizability, We also test different methods on multi-modal mathematical benchmarks including MathVision Wang et al. (2024a) and MathVerse Zhang et al. (2025), which has much lower similarity with our example problem bank. The results are shown in Tab. 2. Problem-level few-shot learning not only fails to enhance reasoning performance but can also have a negative impact, while our method continues to achieve appreciable improvements, demonstrating better general applicability.
Robustness We also manually decrease the similarity between the examples and the problems by selecting the t_th similar example during reasoning to evaluate the robustness. The result is shown in Tab. 3. We can observe that traditional problem-level in-context learning suffers from a severe decrease and is even worse than 0-shot learning when t is larger than 4. In contrast, our method does not show a significant decline and is consistently better than the 0-shot reasoning.
@@ 4.3 Construction of Example Problem Bank
To better align with the steps in reasoning, we propose constructing a step-level problem bank based on the reasoning content rather than grammatical divisions. To prove our assumption, we compare our approach with a commonly used strategy that constructs steps based on grammatical segmentation, using periods ’.’ as the delimiter, on the same dataset PRM800K and under identical conditions. Results are presented in Tab. 5. Our method largely outperforms those using periods as a delimiter.
@@ 4.4 Comparison of Retrieving Strategies
The key factor of in-context learning lies in the relevance of the retrieved examples. At the finer-grained step level, designing an appropriate retrieval strategy becomes even more crucial and challenging. Therefore, we propose the first-try strategy, which involves understanding what the model currently needs to reason about using a first attempt and then searching the problem set for similar steps to guide the model in fully outputting the current step. To validate the effectiveness of this method, we compare it with several other strategies mentioned in Sec.3.3, retrieving by the entire reasoning path $s_{i-1},s_{i-2},\ldots,s_{1},q$ or only by the immediately preceding step $s_{i-1}$.
@@
Table 6: Comparison of different retrieval strategies in step-level in-context learning. ’Path’ represents retrieving by the reasoning path including all previous steps $s_{i-1},s_{i-2},\ldots,s_{1}$ and question $q$, while ’Pre-Step’ represents retrieving by only the immediately preceding step $s_{i-1}$. The best results are in bold.
@@ 4.5 Example-guided Step-level Tree Search
Reason Verify AMC12 AMC10 MATH Avg
w/o tree-search 53.6 55.8 73.4 60.9
58.7 59.0 77.8 65.2 (+4.3)
64.4 62.2 79.2 68.6 (+7.7)
61.6 60.4 78.2 66.7 (+5.8)
65.2 63.6 79.4 69.4 (+8.5)
The reasoning capability of the reason model and the verifying capability of the critic model are two core factors of step-level tree search methods, and our strategy can bring benefits in both ways. On one hand, it can improve the accuracy of generating candidate nodes using the previously mentioned first-try strategy when reasoning nodes are generated. On the other hand, it can increase the accuracy of evaluation by introducing similar examples during critic model assessments and therefore ensures that the correct reasoning nodes are more likely to be preserved. These can be decoupled, allowing us to demonstrate the effectiveness of each component through ablation studies.
We utilize GPT-4o as the reason model, GPT-4o-mini as the PRM and adopt the Pairwise Preference Reward Model (PPRM) configuration Zhang et al. (2024b) to ensure a more robust evaluation. Detailed settings will be listed in the appendix.
Tab. 7 presents the results of integrating in-context learning into the reasoning and evaluation phases of Tree Search methods. The results of this ablation study indicate that introducing example steps can enhance both the reasoning and verifying capabilities of tree search methods. Therefore both approaches contribute to the improvement of overall reasoning performance.
@@ 5 Conclusion
We propose BoostStep, providing fine-grained guidance during the reasoning process by searching for similar steps from a step-level example problem bank according to the first-try reasoning attempt. BosotStep is a strong and general approach, enhancing the model’s reasoning capabilities and reducing the dependency on the similarity of the example problem set. It demonstrates better performance, potential, generalizability and robustness comparing to traditional problem-level few-shot learning. Moreover, our method can also enhance the reasoning and evaluation capability of step-level tree search methods by introducing similar steps in reasoning and verifying phases.
6 Limitations
Currently, our example problem bank is entirely sourced from PRM800k, resulting in a relatively homogeneous distribution of example problems and example steps. Although our method has more potential to guide more difficult problems with much simpler examples, a greater quantity and more diverse distribution of example problems can evidently provide more effective guidance for addressing a range of problems.
Furthermore, the TF-IDF retriever used is based on modeling language term frequency directly and thus lacks an understanding of mathematical content, which limits its retrieval capabilities on math problems. Utilizing a retriever specifically designed for mathematical problems can certainly enhance the quality of retrieval.
@@ A.3 Benchmarks
For multi-modal math benchmarks, we use MathVision-Mini Wang et al. (2024a) and vision-dominant version of problems in MathVerse-Mini  Zhang et al. (2025).
Appendix B Detailed Setup for Example-Guided Step-Level Tree Search
In the setup for tree search methods, we utilize GPT-4o as the reason model and employ GPT-4o-mini as the Process-supervised Reward Model (PRM). For the PRM, we adopted the Pairwise Preference Reward Model (PPRM) configuration Zhang et al. (2024b). Specifically, PPRM transforms the absolute rewards calculation into preference predictions between solutions to calculate rewards. This approach reduces the variability associated with scoring characteristics and thus leads to a more robust and consistent evaluation of different solutions.